If the mass of a sphere with radius R is M then, what will the mass of the sphere with radius R/2 ?

Grade:11

2 Answers

Khimraj
5 years ago
mass = density/volume as density constant and volume is
Khimraj
5 years ago
 
mass = density/volume
as density constant and
volume is is proportional to cube of radius
M1/M2 = (R1/R2)3
M/M2 = (R/(R/2))3
M2 = M/8